OpenStudy (anonymous):

An equation is shown below: 1 over 2 multiplied by x plus 3 over 2 equals 2 to the power of x What is the solution to the equation? x = 1 x = 2 x = 4 x = 8

OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x − 2y = −4 3x + y = 12 A.) −2y − 4 B.) 2y − 4 C.) 2y + 4 D.) −2y + 4

OpenStudy (anonymous):

idk bout these

pooja195 (pooja195):

\[\huge~\rm~\bf~\frac{ 1 }{ 2 }\times x+\frac{ 3 }{ 2 }=2^x\]

OpenStudy (anonymous):

@pooja195 ah ok. so would A be correct for the first one

OpenStudy (anonymous):

and would D be correct for the second question?

OpenStudy (anonymous):

@pooja195 hello?

OpenStudy (dakid88):

A

OpenStudy (tkhunny):

Let's see, \(\dfrac{1}{2}x + \dfrac{3}{2}\) will NOT be an integer unless x is odd. You should see why. \(2^{x}\) will always be in integer if x is an integer. The only answers given are integers. The only POSSIBLE answer is x = 1. All the other choices are even and won't work.

OpenStudy (anonymous):

@tkhunny thank you thats what i was thinking.. do you know about the second one?

OpenStudy (anonymous):

i think D

OpenStudy (tkhunny):

−x − 2y = −4 Isolate x x = 4 - 2y If you were thinking that other argument, you should not be having trouble with this. :-(

OpenStudy (anonymous):

@tkhunny wait not D?

OpenStudy (anonymous):

Wait! haha its D.. thank you! (: